Minutes of the 25th SOHO SWT Meeting

Rutherford Appleton Laboratory, Oxfordshire, UK 27–29 April 1998 Contents

1

1 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Vicente Domingo welcomed all participants and informed the SWT that, as of July 1, he and Art Poland will step down as Project Scientists. Their successors are J. Gurman for NASA and B. Fleck for ESA.

Actions revision

Action 24-1: closed on MDI to reply to the proposed roll steering law update scheme. Input to PS before End of October.

Action 24-2: closed on EIT to submit by e-mail detailed proposal for S/C off-pointing.

Action 24-3: closed on Phil Judge to provide formal JOP on “SOHO Coronal Observations in Support of the Feb 26 1998 Eclipse”. (already closed on 23 October)

2 Status of Experiments

— see annex 3

3 Status of Archives

L. Sanchez, I. Scholl and D. Giaretta gave an update on the status of the SOHO archive at GSFC and the efforts at MEDOC and RAL (see annex 4). They also gave a demonstration of the new user interface for the SOHO archive, which is being developed at MEDOC.

4 Spacecraft Status

Presentation^{1} by H. Schweitzer (see annex 5).

5 Instrument Performance, Status and Trend Analysis

Presentation^{1} by V. Domingo (see annex 6).

6 Science Operations and Data Status

Presentation^{1} by L. Sanchez (see annex 7).

7 ISTP/SOHO Re-engineering Status

R. Mahmot and H. Benefield repeated^{1} the key elements of their presentation on the ISTP/SOHO re-engineering status from the previous day (annex 8).

1given in a technical briefing for the SWT and D/Sci as part of the SOHO Mission Extension Celebration on Tuesday morning prior to the press conference

8 TRACE Update

A. Title gave a status update of the TRACE mission and presented some of TRACE’s spectacular first light images and movies. Some of them are available on the Web at: http://vestige.lmsal.com/TRACE/

9 Science Priorities for the Next 6 Months

There was broad agreement that the next couple of months will be dominated by collaborations with TRACE, at least for CDS, MDI and EIT. In August: 2nd Whole Sun Month campaign.

10 Future S/C Maneuvers and Operations

H. Benefield presented the original plan for the roll/off-pointing maneuver end of June:

• 6/23: Gyro calibration / Momentum management

• 6/25: Yaw offpointing(40 arcmin east)

• 6/26–6/27: 360^{∘} roll maneuver (24 hr duration)

• 6/28: Momentum management / station keeping

• 6/29: SSU software patch

After some lively discussion it was agreed that the pointing disturbances should be clumped together as much as practical, in order to limit the disturbances for the helioseismology instruments as much as possible. GOLF and VIRGO in particular expressed severe concerns that their data will be compromised due to thermal disturbances. In order to allow shorten the whole maneuver activites by one day, CDS offered to waive their request for a 24 h waiting time between the momentum management and the off-pointing. As such, CDS would not be involved in the off-pointing observations. Some members of the EIT team felt that an off-pointing of 40’ in one direction would be preferable to two 25’ off-pointings in both East and West. H. Schweitzer confirmed that the slew to 40’ takes about 40 minutes, so that the whole off-pointing maneuver could be done in less than 2 hours (with an half hour dwell). The agreed plan for the roll/off-pointing maneuver end of June reads:

• 6/24: Gyro calibration / Momentum management

• 6/25: Yaw offpointing(40 arcmin west)

• 6/26–6/27: 360^{∘} roll maneuver (24 hr duration)

• 6/28: Momentum management / station keeping

• 6/29: SSU software patch

It was furthermore agreed to perform a 10’ offpointing to the North for a TRACE collaboration (coronal hole/polar plume campaign). Duration: 8 hours. Target date: 3 September 1998.

Next station keeping maneuver: end October / early November. No roll/off-pointing activities planned for that.

Long DSN contacts:

1 May - 6 May 12 June - 17 June 13 July - 17 July 12 Aug - 17 Aug 4 Sep - 9 Sep

11 MEDOC Campaigns

J.-C. Vial presented plans for future MEDOC campaigns (see annex 9). His proposal to have MEDOC campaign #3 from 8 to 21 June was not accepted by the SWT. CDS, UVCS and TRACE in particular expressed their reluctance to shift the operations from the EOF to MEDOC during that time, for various reasons (too short notice, still within 30-day priority plan of TRACE, certain planners already assigned to the EOF, etc.). As MEDOC personnel could not support a campaign in July or August, it was decided to drop it altogether. The following dates were agreed for MEDOC campaigns # 3-5:

• #3: 7–27 September 1998

• #4: 11–31 January 1999

• #5: 2–15 August 1999 (during eclipse)

12 Future Public Outreach Activities

B. Fleck introduced Jean-Francois Hait to the SWT. J.-F. Hait is a freelance science journalist who was contracted by ESA to help popularize SOHO science results, particularly in Europe.

The following four items were briefly discussed:

• SOHO Results Poster Should have been available for the SOHO extended mission celebrations at RAL. Unfortunately, the resolution and color of the poster front in the final proofs from the film negatives were not satisfactory . In addition, some minor points on the back side needed correction. The corrected version should be available end of May.

• AGU/SPD There will be two SOHO press releases at the AGU/SPD meeting:

1. Solar flare leaves Sun quaking (MDI results)

2. Solar “cat scan” provides most detailed look to date at corona (results from the “Whole Sun Month” campaign)

A. Title suggested to ship thousands of poster to science museums in Boston and distribute them for free. To attract a large croud, this should be widely advertized in newspapers. J. Kohl offered to look into this matter with the press officer from SAO.

A. Title stressed the paramount importance of reaching the broad masses, not just colleagues in the wider scientific community (“mass marketing”).

One of the prime tasks of the ESA science journalist will be to help bring SOHO into science museums in Europe, such as the Cité des Sciences et de L’Industrie, La Villette, Paris, France, or the German Museum in Munich, Germany. Both permanent as well as temporary exhibitions are considered.

• Future Space Science Updates (SSUs)

There are vague plans to have another SSU on MDI results later this year. Topic: 3-D tomography of action regions (birth, sub-surface flows etc.).

• 1999 Eclipse This is a unique chance to bring solar physics to the attention of the broad masses, in particular in the European countries which lie in the path of totality. Preparations for some major outreach activities should be started soon.

13 Meetings and Workshops

B. Fleck presented the list of meetings from the SOHO web pages (see annex 10).

J. Kohl presented the plans and status of preparations for SOHO-7 (see annex 10).

J.-C. Vial offered to host SOHO-8 at Orsay in 1999. Topic: Hot and cool plasma in the solar atmosphere. Target date: week 25 (21-25 June 1999).

The project scientists will follow up plans for a joint workshop with ACE and Ulysses.

14 AOB

SOHO Bibliography Database B. Fleck reminded people to help keep the SOHO bibliography database up-to-date. For a graph of the present content of the database for the various instruments see annex 11. If there are problems using the database input and/or edit facilities, simply send an e-mail with a list of publications to bfleck@esa.nascom.nasa.gov.

New SWT Meeting Cycle A poll among the PIs brought the following result: 6 votes in favour of 3 meetings per year, 2 abstentions, and 4 votes in favour of 2 meetings per year. It was therefore agreed to switch to a cycle of 3 SWT meetings per year: one in February, one in June, and one in October.

Next SWT meetings:

• 15/16 October 1998 at GSFC

• February 1999 at GSFC or ESTEC

• end of June 1999 (6/25?) in Orsay (in conjunction with SOHO-8)

Annex 1: Agenda

7

Annex 2: List of Participants

12

Annex 3: Instruments Status

GOLF (A. Gabriel)

VIRGO (C. Froehlich)

18

MDI (R. Bush)

31

SUMER (P. Lemaire)

47

CDS (R. Harrison)

52

EIT (J.-P. Delaboudiniere)

54

UVCS (J. Kohl)

56

LASCO (R. Howard)

66

SWAN (R. Lallement)

79

CELIAS (P. Bochsler)

82

COSTEP (H. Kunow)

85

ERNE (J. Torsti)

91

Annex 4: Status of SOHO Archives (I. Scholl & D. Giaretta)

100

Annex 5: Spacecraft Status (H. Schweitzer)

105

Annex 6: Instrument Performance, Status and Trend Analysis (V. Domingo)

Annex 7: Science Operations and Data Status (L. Sanchez)

125

Annex 8: ISTP/SOHO Re-engineering Status Report (R. Mahmot & H. Benefield)

Annex 9: MEDOC Campaigns (J.-C. Vial)

147

Annex 10: Future Meetings and Workshops

153

Annex 10: Content of SOHO Bibliography Database

158

Annex 12: Agenda SWT-25 Science Workshop

160